OpenStudy (anonymous):

log 2x + log x = 2 Help please!

geerky42 (geerky42):

HINT: Use this log rule: \(\log(mn) = \log(m) + \log(n)\)

OpenStudy (anonymous):

I don't get that?

geerky42 (geerky42):

You know few logarithm properties, right? Using this rule \(\log(mn) = \log(m) + \log(n)\) would make your problem easier. Using this rule, we have \(\log(2x) + \log(x)~~\Rightarrow~~\log(~(2x)~(x)~)\) Make sense?

OpenStudy (anonymous):

is there a way i can enter it into a calculator? i take notes religiously but this is not in them, this is just confusing me.

geerky42 (geerky42):

Really? Can you tell me what is on your note? Anything you think would be helpful

OpenStudy (anonymous):

Nevermind, i have to take a test now. thank you for your help. (i'm with my hospital homebound teach and we have to stay on a time schedule, i ran out of time)

geerky42 (geerky42):

Okay. Good luck on your test.